OpenStudy (anonymous):

solve (sin^2 x)(cos3x)+(cos3x)(cos^2 x)=1/2 for x over [0,2pi)

OpenStudy (loser66):

factor cos 3x out. you can see the solution.

OpenStudy (anonymous):

ok but how do i do the inverse thing?

OpenStudy (anonymous):

cos = 1/2 at pi/3 right? so would that be the answer?

OpenStudy (loser66):

cos (3x) = 1/2 . For convenient, let 3x = t , so cos t = 1/2 in the interval [0,2pi] you have t = pi/3 and t = 5pi/3 now replace t = 3x so, \[3x = \frac{\pi}{3}\rightarrow x=\frac{\pi}{9}\]do the same with 5pi/3

OpenStudy (anonymous):

oh ok thank you so much!11

OpenStudy (loser66):

got me? good

OpenStudy (anonymous):

\[y=\frac{ \sqrt{x ^{4}+6x ^{2}+9}}{2x ^{2}-10 }\]

OpenStudy (loser66):

solve it?

OpenStudy (anonymous):

do you know how to find all of the asymptotes?

OpenStudy (loser66):

do you know what the asymptotes mean?

OpenStudy (anonymous):

i know what it looks like in a graph, i just don't know how to solve and find them

OpenStudy (loser66):



OpenStudy (anonymous):

x=0 and y=0

OpenStudy (loser66):

good, with x =0, the function undefined, right? with y =0, the same. so, with the fraction as yours, when it's undefined? is not that when the denominator =0?

OpenStudy (anonymous):

oh ok, so if i set the denominator = 0 then i can find the asymptotes

OpenStudy (loser66):

now, go backward, when denominator =0, the fraction is undefined and the value make it undefined is its asymptotes. YOU GOT IT

OpenStudy (loser66):

wud wud, happy when you got it quickly.

OpenStudy (loser66):

ha!! you see , my sentence sounds like a poem, right? hahahaha. I am day dreaming.

OpenStudy (anonymous):

haha yay! ok is that all there is? do i have to do anything different if there is a removable discontinuity or is it any different because there is a square root in the numerator?

OpenStudy (loser66):

just consider the denominator if the question is about asymptote. the numerator is never negative for all x

OpenStudy (anonymous):

ok thanks!!!

OpenStudy (loser66):

yw